# A Spatio-Temporal Aligned SUNet Model FOR LOW-LIGHT VIDEO ENHANCEMENT

### v2

@@ 1 Introduction
Fig. 1: Distortions in cropped images of the ‘Faces2’ sequence. (Left) Normal light. (Middle) Enhanced low light (10% brightness) using histogram matching to the normal light to visualise distortions under low-light conditions. (Right) Normal light plus Gaussian noise. Please zoom-in to see a clear effect of noise.
Fig. 1: Distortions in cropped images of the ‘Faces2’ sequence. (Left) Normal light. (Middle) Enhanced low light (10% brightness) using histogram matching to the normal light to visualise distortions under low-light conditions. (Right) Normal light plus Gaussian noise. Please zoom-in to see a clear effect of noise.
@@
Deep learning techniques have evolved as the most effective tools for image and video processing, in particular the enhancement of low-light content [2, 3, 4]. However, the common approach of sequentially applying image-based techniques to each frame in a video often creates temporal inconsistencies and flickering artefacts. This can, in part, be addressed by enhancing quality on a frame-by-frame basis through temporal smoothing [5]. A further issue that impacts low-light video enhancement is the amount of training data available and it’s quality and diversity. This is compounded by a lack of accurate ground truth information for imagery acquired in low-light conditions, since gathering clean ground truth with accurate brightness, colour and motion is practically impossible. Without spatio-tempoaral alignment between low-light inputs and their ground truth, it is challenging to accurately evaluate the performance of trained models.
Some methods overcome this limitation by creating synthetic datasets or adopting self-supervised learning methods [6]. However, the generalisation on real-world data is often compromised when training on synthetic data since augmenting the dataset with over-simplified synthetic noise, often modelled as zero-mean Gaussian noise, fails to achieve effective results on real world data captured in the wild. This is particularly evident in low-light scenarios where noise characteristics are much more complex than Gaussian noise, as shown in Figure 2. This figure shows a cropped book cover with a human eye, taken from a sequence named ‘Faces2’ with fast linear motion in the BVI dataset [7]. When histogram matching is employed on the low light (10% for noise visualisation), it is noticeable that the shape of the eyeball is slightly shifted and distorted compared to the normal light ground truth. Please zoom-in to better visualise the effect. This is a typical scenario caused by motion blur. The human eye also experiences significant colour flattening and distortion under low-light conditions. Low-light imagery typically loses information during acquisition, including edge distortions, texture loss, motion blur and color changes; these often lead to bias and clipping problems when used to train machine learning methods. Self-supervised learning techniques, although independent of paired input-ground truth datasets and capable of enhancing contrast and brightness by learning subtle content priors, often demand excessive computational complexity. Challenges also arise in the context of color restoration and correction in low-light images where the limited availability of colour information adds to the complexity.
@@ 3.1 Feature alignment
Multiple input images, denoted by $I_{t+i},i\in\{-N,-N+1,\ldots,N-1,N\}$, initially proceed to the feature space via convolutional layers before passing through the alignment module. $N$ represents the number of neighbouring frames at time $t$. This module aligns features of $N$ neighbouring frames with the target frame $I_{t}$ using deformable convolution. This aims to utilise information from adjacent frames in a video while minimising disparities. Pyramid processing and cascading refinement gradually reduce differences between adjacent frames and the target frame. At a lower level $L$, features of two neighbouring frames $I_{t+i}$ and $I_{t}$ concatenate, obtaining the offset $\delta^{L}_{t+i}$ after consecutive convolutional layers, denoted as $L_{a}$ and $L_{b}$. This offset, combined with adjacent frame features $I^{L}_{t+i}$, inputs into deformable convolution, denoted as DC, at that level. They are then upsampled together with a factor of 2 ($C^{L+1}$, where $C$ represents $\delta$ and $I$) and input into the upper layer ($L$+1). For subsequent layers, the offset for each layer concatenates with the upsampled feature from the previous layer. The offset-feature prediction process can be represented as follows:
@@
This iterative 3-layered pyramid process continues until aligned features cascade. A final deformable convolution produces the output of this alignment module - i.e. aligned features. These features are subsequently upsampled in the first up-sampling layer, as depicted in Figure 3, using the Pixel Shuffle technique [25]. This enables patch partitioning in the feature extraction process.

### v3

@@ 1 Introduction
Fig. 1: Distortions in cropped images of the ‘Faces2’ sequence. (Left) Normal light. (Middle) Enhanced low light (10% brightness) using histogram matching to the normal light to visualise distortions under low-light conditions. (Right) Normal light plus Gaussian noise.
@@
Deep learning techniques have evolved as the most effective tools for image and video processing, in particular the enhancement of low-light content [2, 3, 4]. However, the common approach of sequentially applying image-based techniques to each frame in a video often creates temporal inconsistencies and flickering artefacts. This can, in part, be addressed by enhancing quality on a frame-by-frame basis through temporal smoothing [5]. Existing methods often use optical flow for alignment but struggle with accurate motion compensation due to occlusion. A further issue that impacts low-light video enhancement is the amount of training data available and its quality and diversity. This is compounded by a lack of accurate ground truth information for imagery acquired in low-light conditions, since gathering clean ground truth with accurate brightness, colour and motion is practically impossible. Without spatio-tempoaral alignment between low-light inputs and their ground truth, it is challenging to accurately evaluate the performance of trained models.
Some methods overcome this limitation by creating synthetic datasets or adopting self-supervised learning methods [6]. However, the generalisation on real-world data is often compromised when training on synthetic data since augmenting the dataset with over-simplified synthetic noise, often modelled as zero-mean Gaussian noise, fails to achieve effective results on real world data captured in the wild. This is particularly evident in low-light scenarios where noise characteristics are much more complex than Gaussian noise, as shown in Figure 2. This figure shows a cropped book cover with a human eye, taken from a sequence named ‘Faces2’ with fast linear motion in the BVI dataset [7]. When histogram matching is employed on the low light (10% for noise visualisation), it is noticeable that the shape of the eyeball is slightly shifted and distorted compared to the normal light ground truth. This is a typical scenario caused by motion blur. The human eye also experiences significant colour flattening and distortion under low-light conditions. Low-light imagery typically loses information during acquisition, including edge distortions, texture loss, motion blur and color changes; these often lead to bias and clipping problems when used to train machine learning methods. Self-supervised learning techniques, although independent of paired input-ground truth datasets and capable of enhancing contrast and brightness by learning subtle content priors, often demand excessive computational complexity. Challenges also arise in the context of color restoration and correction in low-light images where the limited availability of colour information adds to the complexity.
@@ 3.1 Feature alignment
Multiple input images, denoted by $I_{t+i},i\in\{-N,-N+1,\ldots,N-1,N\}$, initially proceed to the feature space via convolutional layers before passing through the alignment module. $N$ represents the number of neighbouring frames at time $t$. This module aligns features of $N$ neighbouring frames with the target frame $I_{t}$ using a three-level pyramid through deformable convolutions. This aims to utilise information from adjacent frames in a video while minimising disparities. The detailed procedure is described below.
[…]
@@
This iterative process spans three layers. A final deformable convolution generates the output of this alignment module, which are aligned features. These aligned features are then upsampled in the first up-sampling layer, as depicted in Figure 3, using the Pixel Shuffle technique [25], facilitating patch partitioning in the enhancement module.